### v2

# Discovering the underlying analytic structure within Standard Model constants using artificial intelligence

Abstract
This paper presents a search for underlying analytic structures among the fundamental parameters of the Standard Model (SM) using symbolic regression and genetic programming. We identify the simplest analytic relationships connecting pairs of these constants and report several notable expressions obtained with relative precision better than 1%. These results may serve as valuable inputs for model builders and artificial intelligence methods aimed at uncovering hidden patterns among the SM constants, or potentially used as building blocks for a deeper underlying law that connects all parameters of the SM through a small set of fundamental constants.
11footnotetext: Preprint: HEP-ANL-197373, June 26, 2025
1 Introduction
The Standard Model (SM) remains highly successful in describing fundamental particles and their interactions, accurately predicting a broad range of experimental results. However, it is still considered a model rather than a complete theory, as it relies on approximately 20 fundamental parameters that must be determined experimentally [1]. This reliance limits its predictive power. One of the most important open questions in the SM is why these constants have the specific values we observe, and whether a deeper theoretical foundation might explain them.
The Large Hadron Collider (LHC) has enabled numerous stringent tests of the SM, yet no evidence of new physics has been observed. Most searches for new physics at the LHC rely heavily on the SM’s predictive power. However, the SM contains numerous adjustable parameters that must be determined experimentally, which limits the interpretability of the measured cross sections. Moreover, the SM offers limited predictive power for many categories of events produced in $pp$ collisions, leaving significant regions of event topology still largely unexplored [2].
Attempts to identify relationships among the parameters of the SM have a long history [3, 4, 5]. However, no convincing evidence has been found to support the existence of such relationships. As a result, these constants are generally regarded as free parameters of the SM, with no apparent underlying connections.
Finding analytic relationships through brute-force searches or by noticing coincidences in numerical values or measurements was characteristic of the early stages of science - before the development of fundamental theories that could explain connections between seemingly unrelated quantities. However, this empirical approach proved to be useful and sometimes led to the discovery of new laws that revealed the underlying reasons for such coincidences. It is not improbable that a similar strategy could be valuable today, especially with the increasing availability of artificial intelligence (AI) algorithms and greater computational power.
In this paper, we explore the possibility of uncovering underlying structures among the fundamental constants of the SM using symbolic regression and genetic programming - a type of evolutionary algorithm inspired by natural selection and rooted in the broader field of AI. We apply this method to investigate whether analytic relationships among SM constants reveal an inherent structure or potentially point to a deeper, unifying law that governs them.
2 Technical details
Genetic programming (GP) is a powerful technique in symbolic regression that evolves analytic expressions directly from data [6]. By representing mathematical expressions as trees and applying evolutionary algorithms, GP explores a wide solution space unconstrained by predefined model structures, enabling the discovery of complex, previously hidden relationships.
In symbolic regression, the system learns an (symbolic) expression from a set of input-output pairs, often many such pairs, and often there are many input values (representing the dependent variables), and a single output (the independent variable). The generated expressions then contain relations with these variables, functions of these variables, as well as constants. It can then be used to evaluate expressions with inputs not seen in the given data set.
The symbolic computations were performed using a program written in the Picat language [7], a logic-based, multi-paradigm programming language that combines features from logic programming, constraint solving, functional programming, planning, and imperative programming.
The program primarily leverages Picat’s logic programming (non-determinism), functional constructs, and imperative features. It is available on the second author’s website [8].
The specific variant of symbolic regression that is used here (constant identification), is different in some important aspects:
  • The data set for a single run/experiment consists of a single input-output pair:
    <1, value of the physical constant>
    where 1 is the input, and the value of the specific physical constant is the output.
  • The algorithm then tries to learn some relation between the constant 1 and value the physical constant, generating an expression. See below for more on the specific genetic programming method.
  • The found expression does not contain variable names as described above for the general symbolic regression approach. In our version of constant identification, everything in the generated expression is either a function or constants (physical constant or an integer).
  • The equation for the physical constants that is shown below in the paper contains the name of the physical constants, but that is handled by post-processing after the symbolic regression part is done.
Some specific features of the program are:
  • The number of expressions in each population is 1000. In the first run, the expressions are generated randomly from:
    • functions: +, --, *, /, $\sqrt{(}x)$, $x^{2}$, $x^{3}$, $x^{4}$, $\exp(x)$, $\log(x)$;
    • the integers from 1 to 10;
    • and the two or three physical constants involved, the output and input value(s).
    The use of only small integers instead of larger integers or random floats is for simplicity and for the interpretability of an expression.
  • The analytic formulas are evaluated for each expression in a population, and the results that are closest to the output value are kept for the next generation. The formulas that are too far are replaced with previous expressions that are combined from other expressions using the standard genetic programming operators crossover and mutation (with some specific probability). New random formulas may also be added (from the same set as above).
  • When an expression is sufficiently close to the value of output $c$ within the allowed error of the output - then this is considered a solution. All solutions found within the allotted number of runs (1000) are reported. If no solution is found within the allotted number of runs, the program continues to run until a solution is found (and then reported), or until the program times out. The timeout for each experiment, based on a given output–input configuration, was set to 60 seconds.
  • An important feature of the program is that it enforces the inclusion of all input value(s) in each expression. Any expression that does not contain the required input value(s) is removed from the population.
The results presented in this paper are based on 3,249 distinct runs, each involving the testing of an output against one or two input constants, as will be discussed later. Many single runs produced multiple solutions. The results were obtained from multiple runs with different random seed values, which takes into account the inherent stochastic nature of GP. Duplicate solutions were removed during the post-processing step. All runs took in total about 15,000 CPU hours with the Intel E5-2650 v4 processor and 252Gb RAM.
2.1 Inputs for the genetic algorithm
A critical aspect of evaluating analytic relationships between SM parameters is the selection of an appropriate input space for the algorithms. Working with SM parameters presents several challenges. First, most of these parameters have physical units, which restricts the applicability of the algorithms due to dimensional analysis constraints — only certain analytic structures are valid when dimensions are taken into account. To enable numerical evaluation, we focus on dimensionless constants. There are arguments [9] suggesting that the laws of physics should be independent of the choice of units or measuring instruments. Dimensionless constants, in particular, are well-suited for probing the deep structure underlying physical laws.
The second challenge for genetic algorithms lies in the wide range of SM values, spanning several orders of magnitude — from the electron mass to the top quark mass, the latter being the heaviest among all SM particles.
Finally, the uncertainties associated with the SM parameters vary significantly and must be taken into account when evaluating the robustness of any discovered relationships. If one variable has very small uncertainties, it offers little advantage if other input variables have large uncertainties, which will dominate the overall error.
One way to address the issue of physical units in the SM is to normalize all particle masses by dividing them by another parameter that has units of mass. However, this approach introduces a degree of arbitrariness in the choice of the reference mass. A possible option for the divider is the Planck mass [1], $m_{\mathcal{P}}=1.22089(6)\times 10^{19}$ GeV. While theoretically appealing, this choice is impractical for our purposes, as the GP algorithm we employ cannot handle values smaller than $10^{-16}$.
To resolve this, we normalize all SM masses by the mass of the $\rho(770)$ meson, 0.77526±0.000230.77526$\pm$ 0.000230.77526 ± 0.00023 GeV [1]. At first glance [^2] We could also use the mass of any fundamental particle for normalization, such as the electron mass. However, that particle would then be excluded from our GP analysis., this might seem like an unconventional choice. However, it is not without justification for the GP algorithm. The $\rho(770)$ meson holds a special place in the context of quantum chromodynamics, plays a key role in hadronic structure, and even electromagnetic interactions. Its mass lies conveniently in the GeV range, which allows us to rescale the SM mass spectrum into a range well-suited for GP.
Finally, using the uncertainty of the $\rho(770)$ meson in error propagation, when dividing the particle masses by its mass, results in a consistent normalization of uncertainties across all mass values. For GP, it is advantageous to work with uncertainties that are approximately uniform across all inputs, as this improves the stability and comparability of the algorithm’s performance.
For example, the mass of the electron, 0.51099895000(15) MeV, has an exceptionally small relative uncertainty of $3\times 10^{-8}\%$, while the Higgs boson mass, 125.20(11)125.20(11)125.20 ( 11 ) GeV [1], has a much larger relative uncertainty of approximately 0.09%. This disparity in precision poses a challenge for numerical algorithms. However, scaling all masses by the $\rho(770)$ meson mass and propagating its uncertainty to the resulting ratios helps mitigate this issue.
Thus, in our GP algorithm, the maximum precision of the input parameters was set to 0.03%. The input includes two dimensionless constants whose precision significantly exceeds this threshold: the mathematical constant $\pi$, and the fine-structure constant $\alpha$ (or its inverse value $\alpha^{-1}$, which will be used in our study). Although the constant $\pi$ is not a parameter of the SM, it frequently appears in the laws of physics. To maintain consistency, we reduce their precision to $\pi=3.14159(1)$ and $\alpha^{-1}=137.036(1)$. Even with this adjustment, they remain 40–100 times more precise than the most precisely known mass parameter — the electron mass normalized by the $\rho(770)$ meson mass. Increasing the precision of $\pi$ and $\alpha^{-1}$ is not expected to affect the results involving particle masses, as the overall uncertainty will still be dominated by uncertainties in the mass values. However, using overly precise constants can introduce numerical instability and make it more difficult for GP to reproduce known relationships. These validation tests will be discussed later.
With all these assumptions, the input parameters for the GP processing are shown in Table 1. Note that we did not include the PMNS mixing angles [1], which are often considered a part of the fundamental constants of SM. They are more relevant for neutrino physics, but neutrino masses are not considered. The measured precision of their values is rather low which can trigger many “noisy” solutions in the evaluation of relations using GP. We also did not include the speed of light ($c$), which has units. However, one should mention that $c$ is indirectly included in the calculation via the fine-structure constant $\alpha$.
Table 1: Dimensionless constants [1] used for the GP algorithm. All particle masses in this table are divided by the mass 0.77526±0.000230.77526$\pm$ 0.000230.77526 ± 0.00023 GeV of the $\rho(770)$ meson to improve the value range and for reducing the large variability of uncertainties, making them more uniform for GP. The traditional symbol $m$ was chosen for the rescaled masses to simplify the presentation of the results. We require that the uncertainties for $\pi$ and $\alpha^{-1}$ are at most a factor 100 higher than for best measured mass. Making them more precise should not contribute to the results. The table presents both the absolute measured uncertainties (±$\pm\varepsilon$) and the relative uncertainties (±$(\pm\varepsilon^{rel}$, in percent).
Constant Name Value ±$\pm\varepsilon$$\pm\varepsilon^{rel}$ (%)
PI $\pi$ 3.14159 1e-05 0.0003
Fine-structure constant (inv) $\alpha^{-1}$ 137.036 0.001 0.0007
Strong coupling constant at $Z^{0}$ $\alpha_{S}$ 0.1180 0.0009 0.7627
CKM 12-mixing angle $\theta_{12}$ 0.22501 0.00068 0.3022
CKM 23-mixing angle $\theta_{23}$ 0.04183 0.00079 1.8886
CKM 13-mixing angle $\theta_{13}$ 0.003732 9e-05 2.4116
CKM CP-violating Phase $\delta$ 1.147 0.026 2.2668
electron mass $m_{e}$ 0.0006591 2e-07 0.0303
muon mass $m_{\mu}$ 0.1363 4e-05 0.0293
tau mass $m_{\tau}$ 2.292 0.0007 0.0305
u-quark mass $m_{u}$ 0.002786 9e-05 3.2304
d-quark mass $m_{d}$ 0.006062 9e-05 1.4847
s-quark mass $m_{s}$ 0.1206 0.001 0.8292
c-quark mass $m_{c}$ 1.642 0.0059 0.3593
b-quark mass $m_{b}$ 5.3956 0.0092 0.1705
t-quark mass $m_{t}$ 222.583 0.405 0.1820
Z-boson mass $m_{Z}$ 117.622 0.035 0.0298
W-boson mass $m_{W}$ 103.667 0.035 0.0338
H-boson mass $m_{H}$ 161.494 0.149 0.0923
The least precise masses among our inputs are those of the light-flavor quarks, which were taken from the PDG using the $\overline{\text{MS}}$ scheme at a renormalization scale of 2 GeV. The most significant challenge is also when using the CKM mixing angles as inputs. Due to their relatively large uncertainties, there is a considerable risk of generating spurious equations.
It should be emphasized that the presence of the mass of $\rho(770)$ (or any other particle mass used for obtaining dimensionless SM constants) should be considered a handy method for deriving short relationships in GP. It is an auxiliary parameter. The final theory can be free of this constant after successful variable substitutions, which should be performed after obtaining the analytic expressions.
2.2 Search strategy
The input parameters shown in Table 1 will be used for the following studies as described below.
First, this paper aims to identify the simplest analytic relationships of the form $f(a)=b$ and $F(a,b)=c$. By “simplest”, we refer to functions $f$ and $F$ with minimal structural complexity, or, in most cases, the shortest possible expression length. We assume that the laws of nature tend to manifest through the simplest possible analytic relationships.
When searching for analytic dependence between constants, one may encounter very complex solutions. To manage this complexity, we have organized all functions according to their simplicity ranks. The lower the rank, the simpler the solution. To construct the rank for each analytic expression, each mathematical operator was assigned a value as shown in Table 2. These values were summed up to create a rank for the final analytic expression. For example, $\delta=\pi-2$ has a rank of 6 because it includes the constant $\pi$ (contributing 1 to the rank), a plain integer (value 2) and a subtraction operation (contributing 3 to the rank).
Table 2: Assigned values for evaluating ranks for the obtained analytic expressions. These values will be sum up for evaluating the rank of each analytic solution.
Value Mathematical token
1 physics constant (float) and $\pi$
2 plain integer (1…10)
3 addition and subtraction
4 multiplication and division
5 $\sqrt{x}$, $x^{2}$, $x^{3}$ and $x^{4}$
6 $\exp(x)$ and $\log(x)$ ($\log$ base of $e$ of $x$), $\cos(x)$, $\sin(x)$, $\tan(x)$, $\arctan(x)$
Under this agreement for defining analytic ranks, the algorithm prioritizes basic algebraic operations (such as addition, subtraction, multiplication, and division) while assigning larger values to exponential, logarithmic, and trigonometric functions. Note that the analytic ranks depend on the default representations in the Picat output. No attempt is made to simplify expressions.
We will refrain from making additional theoretical assumptions, aiming to keep the solutions as open to potential theories as possible. Even if no apparent structure emerges that links most of the SM parameters, these snippets of relationships may still prove valuable for theory builders, when they are combined together. They could serve as building blocks for discovering more intricate laws that potentially unify some or all of these fundamental constants.
2.3 Validation
As a sanity check, we have validated the genetic programming approach using known relationships. Although these relationships are not discussed in the context of the SM or any theory, they were established in the past, and well suited for our purpose of testing the algorithm.
The first relationship is [10]:
$4\pi^{3}+\pi^{2}+\pi=\alpha^{-1}.$ (2.1)
This expression is particularly useful since both parameters are the most precise in our inputs, thus it provides a reliable test of the GP algorithm. This relation should be true within the precision chosen by us in Table 1.
The second relationship for our benchmarks is [11]:
$\sqrt{m_{Z}\>m_{t}}=m_{H}.$ (2.2)
There is no immediate explanation for why such relationships hold, which may suggest that they are merely coincidental. However, if our calculations are able to reproduce them, we will consider the program to be sufficiently robust.
Our program detected Eqs. 2.1 and 2.2 expressions using special test runs. The obtained simplicity ranks were 30 and 20, respectively. The difference $m_{H}-\sqrt{m_{Z}\>m_{t}}$ = -0.310 is greater than the absolute uncertainty allowed of 0.149 on the Higgs mass shown in Table 1. When GP finds this equation, it is not considered a solution.
3 Results
The expressions with the lowest analytic ranks 6 and 7 are shown in Tables LABEL:tab:long6 and LABEL:tab:long7, respectively. The regularities appear rather trivial, which can be checked by hand. Table LABEL:tab:long10 of the Appendix shows the results with the rank 10. Note that the analytic expressions with ranks 8 and 9 do not exist because of the default construction of the ranking system. We intentionally retained repeated snippets involving the same variables, as they may help model builders identify interesting patterns. In total, about 530,000 analytic relations were obtained [12] up to the rank 80, with and without rescaling by the $\rho(770)$ meson mass.
Table 3: Analytic expressions for the rank 6. The values of $\Delta D$ show the relative difference between the predicted and target values (in percent).
Nr Expression for rank 6 $\Delta D$ (%)
1 $\alpha_{S}=m_{\mu}/\delta$ 0.71
2 $\delta=1+m_{\mu}$ 0.93
3 $\delta=\pi-2$ 0.47
4 $\delta=m_{\mu}/\alpha_{S}$ 0.70
5 $\theta_{23}=\theta_{12}/m_{b}$ 0.30
Table 4: Analytic expressions for the rank 7. The values of $\Delta D$ show the relative difference between the predicted and target values (in percent).
Nr Expression for rank 7 $\Delta D$ (%)
1 $\delta=\exp(m_{\mu})$ 0.08
2 $\delta=\log(\pi)$ 0.20
3 $\delta=m_{\tau}/2$ 0.09
The tables also show the relative difference for each analytic snippet. It is defined as $\Delta D=100\%\times|y-y_{t}|/y_{t}$ (expressed as a percentage), where $y_{t}$ is the target value and $y$ is the predicted value. All relations have $\Delta D<\varepsilon^{rel}$ on the target constants according to the definition of GP. The formulas with $\Delta D$ above 1% are not shown, to reduce the sizes of the tables. This requirement truncates the presentation for some targets, which usually correspond to $u,d$-masses, the CKM mixing angles and CP-violating phase, since they have relative uncertainties $\varepsilon^{rel}$ above 1%, see Table 1. Other target constants should not be affected by the 1% requirement, since their $\Delta D$ values are smaller than $\varepsilon^{rel}$ listed in Table 1.
The results shown in Tables LABEL:tab:long6 - LABEL:tab:long10 are dominated by the expressions for $\delta$, which has the largest absolute and relative uncertainties ($\varepsilon$ and $\varepsilon^{rel}$) among all the input variables. Among all the ranks obtained, leptons had the fewest solutions due to their relatively small mass uncertainty. All relationships from the symbolic regression have been preserved, even if they may appear highly unlikely. We remind that this choice reflects our intent to remain as theory-agnostic as possible, in case these analytic snippets are combined and simplified, and later used to inform or construct a model that connects all SM constants.
Most of the obtained GP solutions, if not all, are likely coincidental and represent numerical noise. Identifying meaningful dynamic relationships based on symmetry or other mathematical and physical principles remains an important direction for future work. A visual inspection of the lowest-rank expressions revealed no notable patterns, such as systematically repeated constants or functions, that might indicate underlying regularities. The frequencies of integers from 1 to 10 appear to be uniformly distributed, with approximately the same mean value.
Approximately 10% of the obtained relations satisfy the dimensional analysis if the rescaled masses listed in Table 1 are replaced with their original values expressed in physical units (MeV or GeV). Note that this replacement presumes the same numerical precision as that used for the rescaled masses in the GP inputs; otherwise, the expressions may not satisfy the GP output precision criteria.
A preliminary visual inspection of the lowest-rank expressions reveals relationships that appear to be both remarkably simple and precise, warranting further investigation. For example, Table LABEL:tab:long10 has an expression $m_{t}=m_{W}+m_{W}\>\delta$ connecting the top-quark mass $m_{t}$ with the $W$ mass and $\delta$. The rank 15 snippets [^3] Tables with analytic ranks above 10 are not shown to reduce the size of the paper. contain another relation $m_{H}=m_{t}-(m_{b}+m_{t}/4)$ connecting the Higgs mass $m_{H}$ with $m_{t}$ and the $b-$quark mass. Another relation of the same rank is $m_{u}=2m_{e}/\sqrt{\theta_{12}}$, where $m_{u}$ is the mass of the $u$-quark, $m_{e}$ is the mass of the electron and $\theta_{12}$ is the CKM 12 mixing angle. The found expression $m_{H}=5\>m_{W}-3\>(m_{t}-m_{W})$ connects the Higgs mass with the top-quark mass and the $W$ mass. All such expressions are valid for the original mass units, without scaling by the $\rho(770)$ mass. The precision of these analytic relations significantly exceeds that of Eq. 2.2.
Many relations involving the Higg mass are valid only for the dimensionless constants listed in Table 1. For example, $m_{H}=m_{t}-6\>\sqrt{m_{W}}$ connects $m_{H}$, $m_{W}$ and $m_{t}$. It does not pass the dimensional analysis when using the original masses, i.e. without rescaling the masses by the $\rho(770)$ meson mass. However, the mass of this meson can easily be substituted using Table LABEL:tab:long6 or LABEL:tab:long7, as we will discuss this later.
As an additional test, the GP studies were also conducted on the SM inputs without rescaling the lepton, quark, and heavy boson masses [1] by the $\rho(770)$ mass. In this case, the majority of snippets do not pass dimensional analysis, and the expressions for the electron and muon masses cannot be recovered due to the large precision on such masses. It was determined that the most interesting relations that pass dimensional analysis can also be recovered in the original analysis with the rescaled masses.
One of the known property of the SM is that the masses of fermions, as well as the $m_{Z}$ and $m_{W}$ boson masses, tend toward zero when the Higgs boson mass is set to zero. This behavior arises from their dependence on the Higgs vacuum expectation value through the Yukawa couplings, which are free parameters in the theory [13]. We identified several simplest relations connecting the quark and $Z/W$ masses with the Higgs mass, which have the above mentioning property, and also pass the dimensional analysis. Here are some examples with small analytic ranks:
$m_{u}$ $=\frac{m_{H}}{3\>(\alpha^{-1})^{2}},\quad m_{d}=m_{u}(1+\delta),$ (3.1)
$m_{s}$ $=\frac{m_{d}}{\theta_{12}^{2}},\quad m_{c}=m_{d}\>\left(2+\frac{1}{\theta_{13}}\right),$ (3.2)
$m_{b}$ $=\frac{m_{c}\>\pi^{2}}{3},\quad m_{t}=m_{b}\>\pi(\pi+10),$ (3.3)
$m_{Z}$ $=m_{t}\>\log(\delta)(5-\delta),\quad m_{W}=m_{t}/(1+\delta).$ (3.4)
These relations hold for the masses presented in Table 1, as well as for the original masses of quarks and bosons [1] with the physical units. Analytical expressions for lepton masses as linear functions of the Higgs mass were not found. Nevertheless, these examples illustrate that the reduction of the free parameters of the SM is possible, since 8 masses are expressed via the Higgs mass and 4 measured constants ($\alpha^{-1}$, $\delta$, $\theta_{12}$ and $\theta_{13}$). The origin of these expressions may be purely coincidental, and can only be understood within the framework of a specific model, which is beyond the scope of this paper.
It will be interesting to assess the accuracy of the derived expressions as increasingly precise measurements of the SM parameters become available over time. The published relationships can be used for developing theories or as inputs for various artificial intelligence techniques to discover hidden analytic structures. For example, the analytic snippets can be simplified, combined, variables can be substituted, and a dimensional analysis can be applied after converting the masses back to their original physics units.
We should recall that the introduction of the non-fundamental particle in our study, the $\rho(770)$ meson, was made for convenience, to create the dimensionless inputs for GP. It is a parameter that defines the limits of numerical precision for the GP and, consequently, the applicability of any potential theory connecting the SM constants, should such a theory exist. In principle, any particle can be used for this purpose, as long as the rescaled masses remain within a numerically suitable range for the GP framework. All rescaled masses in higher-rank analytical expressions can be straightforwardly converted to actual masses in units of GeV or MeV using the most suitable relation from Table LABEL:tab:long6 or LABEL:tab:long7, followed by appropriate substitutions. For example, one can use $\delta=\exp(m^{\prime}_{\mu}/m^{\prime}_{\rho})$ from Table LABEL:tab:long7, where $m^{\prime}_{\mu}$ and $m^{\prime}_{\rho}$ denote the muon and $\rho(770)$ meson masses in GeV or MeV. From this expression, the value of $m^{\prime}_{\rho}$ can be determined and subsequently applied across all dimensionless expressions derived in this study. Determining the optimal analytic expression for the $\rho(770)$ mass, however, lies beyond the scope of this paper. This task should be pursued in future analyses of the derived analytic snippets, with the goal of identifying relationships among fundamental constants that enable the construction of physics models with a minimal set of parameters.
4 Conclusion
In this paper, we derived analytic expressions - referred to as “snippets” - connecting the fundamental constants of the SM using an approach based on GP. We presented results up to analytic rank 10, limiting the solutions to a relative precision of at most 1%. The results with a relative difference greater than 1%, corresponding to the $u$- and $d$-quark masses, the CKM mixing angles, and the CP-violating phase, are not included in this article. Analytic expressions with ranks above 10, without any restriction on predicted precision $\Delta D$, and without rescaling by the $\rho(770)$ meson mass, are available [12] from the authors.
A few known (though likely coincidental) relationships were successfully recovered. Surprisingly, the obtained GP results reveal a significant number of simple analytic patterns. Some of them can pass dimensional analysis. However, no evident patterns emerged from the relationships between the constants that could indicate an underlying simple mathematical structure with a limited number of free parameters. Although we expect that most of the expressions, undoubtedly, arise by pure chance, it is possible that some may hint at an underlying structure or deeper theory linking these fundamental constants - should such a theory exist. The coincidental nature of some (or all) of the snippets will be revealed over time as more precise data on the SM parameters become available.
We believe that the compiled lookup library of analytic snippets will be useful for model builders. We also hope that the derived expressions will serve as a valuable resource for future AI-driven investigations, providing building blocks for uncovering a potential underlying law that connects all parameters of the SM through a small set of fundamental constants.
Acknowledgments
We would like to thank Dr. C. K. Zachos for valuable discussions. The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (“Argonne”). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. Argonne National Laboratory’s work was funded by the U.S. Department of Energy, Office of High Energy Physics under contract DE-AC02-06CH11357.
Appendix
Table LABEL:tab:long10 lists the analytic snippets for the rank 10. We retain the mathematical notations used in the Picat language to facilitate future processing. They employ round brackets for functions such as $\sqrt{x}$, while the operators $pow2(x)$, $pow3(x)$, $pow4(x)$ refer a power of 2, 3 and 4, respectively.
Table 5: Analytic expressions for the rank 10. The values of $\Delta D$ show the relative difference between the predicted and target values (in percent).
Nr Expression for rank 10 $\Delta D$ (%)
1 $\alpha_{S}=(m_{Z}-m_{W})/m_{Z}$ 0.54
2 $\alpha_{S}=m_{e}/(m_{u}+m_{u})$ 0.24
3 $\alpha_{S}=m_{s}-pow2(\theta_{23})$ 0.72
4 $\alpha_{S}=m_{s}-pow3(m_{s})$ 0.72
5 $\alpha_{S}=m_{s}-pow3(m_{\mu})$ 0.06
6 $\alpha_{S}=m_{s}-pow4(\theta_{12})$ 0.03
7 $\alpha_{S}=m_{s}/(\delta-m_{s})$ 0.43
8 $\alpha_{S}=m_{\mu}-pow2(m_{\mu})$ 0.24
9 $\delta=(\theta_{12}+m_{c})/m_{c}$ 0.87
10 $\delta=(\theta_{23}+m_{d})/\theta_{23}$ 0.18
11 $\delta=(m_{t}-m_{W})/m_{W}$ 0.009
12 $\delta=1+(m_{\mu}+\theta_{13})$ 0.61
13 $\delta=1+m_{\mu}+m_{e}$ 0.88
14 $\delta=1+m_{\mu}-m_{e}$ 0.99
15 $\delta=2+(m_{\tau}-\pi)$ 0.30
16 $\delta=\alpha^{-1}/\alpha^{-1}+m_{\mu}$ 0.93
17 $\delta=\alpha_{S}+m_{s}/\alpha_{S}$ 0.61
18 $\delta=\pi-(2+\theta_{13})$ 0.80
19 $\delta=\pi-(2+m_{e})$ 0.53
20 $\delta=\pi-m_{u}-2$ 0.71
21 $\delta=\theta_{13}+\pi-2$ 0.15
22 $\delta=m_{b}/m_{b}+m_{\mu}$ 0.93
23 $\delta=m_{c}/m_{c}+m_{\mu}$ 0.93
24 $\delta=m_{d}+(\pi-2)$ 0.06
25 $\delta=m_{d}/m_{d}+m_{\mu}$ 0.93
26 $\delta=m_{e}+\pi-2$ 0.41
27 $\delta=m_{e}/m_{e}+m_{\mu}$ 0.93
28 $\delta=m_{s}+m_{s}/\alpha_{S}$ 0.38
29 $\delta=m_{s}/(\theta_{12}-m_{s})$ 0.70
30 $\delta=m_{t}/m_{t}+m_{\mu}$ 0.93
31 $\delta=m_{u}-(2-\pi)$ 0.23
32 $\delta=m_{u}/m_{u}+m_{\mu}$ 0.93
33 $\delta=m_{\mu}+1+m_{u}$ 0.69
34 $\delta=m_{\mu}+\alpha_{S}/\alpha_{S}$ 0.93
35 $\delta=m_{\mu}+\pi/\pi$ 0.93
36 $\delta=m_{\mu}+\theta_{12}/\theta_{12}$ 0.93
37 $\delta=m_{\mu}+\theta_{13}/\theta_{13}$ 0.93
38 $\delta=m_{\mu}+\theta_{23}/\theta_{23}$ 0.93
39 $\delta=m_{\mu}+m_{H}/m_{H}$ 0.93
40 $\delta=m_{\mu}+m_{W}/m_{W}$ 0.93
41 $\delta=m_{\mu}+m_{Z}/m_{Z}$ 0.93
42 $\delta=m_{\mu}+m_{d}+1$ 0.40
43 $\delta=m_{\mu}+m_{s}/m_{s}$ 0.93
44 $\delta=m_{\mu}+m_{\mu}/m_{\mu}$ 0.93
45 $\delta=m_{\mu}+m_{\tau}/m_{\tau}$ 0.93
46 $\delta=m_{\tau}-m_{\mu}-1$ 0.76
47 $\delta=\sqrt{(}m_{c})-m_{\mu}$ 0.16
48 $\theta_{12}=\sqrt{(}\alpha_{S})-\alpha_{S}$ 0.22
49 $\theta_{13}=m_{d}/(m_{c}-m_{d})$ 0.71
50 $\theta_{23}=m_{s}*(\theta_{12}+m_{s})$ 0.36
51 $\theta_{23}=m_{\mu}/(\pi+m_{\mu})$ 0.59
52 $\theta_{23}=m_{\mu}/(m_{c}+m_{c})$ 0.78
53 $m_{H}=(\theta_{12}+\theta_{12})/m_{u}$ 0.02
54 $m_{W}=\delta+m_{Z}/\delta$ 0.03
55 $m_{W}=m_{H}*m_{c}-m_{H}$ 0.01
56 $m_{Z}=\delta*(m_{W}-\delta)$ 0.03
57 $m_{b}=7+(\theta_{23}-m_{c})$ 0.08
58 $m_{b}=m_{c}*(m_{c}+m_{c})$ 0.06
59 $m_{b}=m_{\mu}+pow2(m_{\tau})$ 0.11
60 $m_{b}=pow4(m_{c}-\alpha_{S})$ 0.02
61 $m_{c}=(m_{H}+m_{W})/m_{H}$ 0.005
62 $m_{c}=2-(m_{\mu}+\theta_{12})$ 0.20
63 $m_{c}=7+(\theta_{23}-m_{b})$ 0.26
64 $m_{c}=\delta+\delta/m_{\tau}$ 0.33
65 $m_{c}=\theta_{12}/m_{s}-\theta_{12}$ 0.08
66 $m_{c}=pow2(m_{\mu}+\delta)$ 0.30
67 $m_{c}=pow3(\delta)+m_{\mu}$ 0.20
68 $m_{c}=\sqrt{(}\pi)-m_{\mu}$ 0.36
69 $m_{d}=\theta_{13}*(m_{c}-\theta_{13})$ 0.86
70 $m_{d}=\theta_{13}/m_{c}+\theta_{13}$ 0.94
71 $m_{s}=\alpha_{S}*(\delta-\alpha_{S})$ 0.68
72 $m_{s}=\alpha_{S}+pow2(\theta_{23})$ 0.70
73 $m_{s}=\alpha_{S}+pow3(m_{\mu})$ 0.06
74 $m_{s}=\theta_{12}/(\theta_{12}+m_{c})$ 0.07
75 $m_{s}=\theta_{23}+\sqrt{(}m_{d})$ 0.76
76 $m_{s}=m_{\mu}-\alpha_{S}*m_{\mu}$ 0.32
77 $m_{s}=pow3(\alpha_{S})+\alpha_{S}$ 0.79
78 $m_{s}=pow3(m_{c}-\delta)$ 0.57
79 $m_{s}=pow4(\theta_{12})+\alpha_{S}$ 0.03
80 $m_{t}=1+(m_{Z}+m_{W})$ 0.13
81 $m_{t}=m_{W}+\delta*m_{W}$ 0.004
82 $m_{t}=m_{c}*(\alpha^{-1}-m_{c})$ 0.12
83 $m_{u}=(\theta_{12}+\theta_{12})/m_{H}$ 0.02
84 $m_{u}=m_{e}/(\alpha_{S}+\alpha_{S})$ 0.24